OpenStudy (anonymous):

–228 + 384 + (–189) I am in a rush.

OpenStudy (amonoconnor):

-228 + 384 + (-189) = (-228) + 384 - 189 *Doing any subtraction first (before addition that is) is a dependable way to be accurate with order of operations in math. So... (-228) + 384 - 189 = (-228) + (384 - 189) = (-228) + (195) = -33 Make sense? :)

OpenStudy (amonoconnor):

Does that make sense @Bella5 ?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (amonoconnor):

You bet! medal perhaps? ;)

OpenStudy (anonymous):

OK :)

OpenStudy (amonoconnor):

;)